1.Goreliebium 2ooo has a half-life of 4 weeks.If there are 100 units of goreliebium 2000 in the begining,
a.how much will be left in 3 weeks?
b.how much will be left in 7 weeks?
2. The average height of a longleaf pine as a function of its age is given by
h(t)=120(1-e^-.o4t). How tall is a typical tree after 30 years?After 100 yrs.?After 150 yrs.?
not from my math 161 book. instructor came up with same.

Answer by scott8148(6628)

half-life means that half of the original material will be gone in the stated time

the formula is a=A((1/2)^(t/h)) where a is the amount remaining, A is the starting amount, t is the time , h is the half-life

a) a=100(.5^(3/4))=100(.595)=59.5

b) a=100(.5^(7/4))=100(.297)=29.7

2)a) h=120(1-e^((-.04)30))=120(1-e^(-1.2))=120(1-.301)=83.9

2)b) h=120(1-e^((-.04)100))=120(1-e^(-4))=120(1-.0183)=117.8

2)c) h=120(1-e^((-.04)150)=120(1-e^(-6))=120(1-.00248)=119.7